# Configurations, Tessellations and Tone Networks

### v4

@@ II The Tonnetz as a Levi Graph
An $L$-transformation sends a major chord notch clockwise on the Levi graph to the next minor chord, and it sends a minor chord one step counterclockwise to the previous major chord. An $R$-transformation sends a minor chord notch clockwise on the Levi graph to the next major chord, and it sends a major chord one step counterclockwise to the previous minor chord. A $P$-transformation sends any chord across the Levi graph to its mode reverse.
The general pattern of a $2p$-hexacycle is as follows. Beginning at the chord $CM$ the cycle moves two steps clockwise along the graph to $FM$; then it crosses the graph with a $P$-transformation to $Fm$; then it moves two steps counter-clockwise to $Cm$, and finally back to $CM$ with another $P$-transformation. From left to right the string of operations applied to the initial chord $CM$ in this example going clockwise thus takes the form $RLPRLP$, or else $PLRPLR$ going counterclockwise. The other remaining eleven $2p$-hexacycles can be obtained by moving the initial point around in a circle of fourths, that is to say, from $CM$ to $FM$ to $B_{\flat}M$ and so on. Each tone is contained in three distinct $2p$-hexacycles. The $3p$-hexacycle (‘blunted triangle’) beginning at $CM$ is generated by the string $PLPLPL$ clockwise and the string $LPLPLP$ counterclockwise. This is the unique hexatonic cycle containing that chord. The $1p$-octacycle (‘octabeanie’) beginning at $CM$ is generated by the string $RLRLRLRP$. And the $4p$-octacycle (‘four-cornered hat’) beginning at $CM$ is generated by $RPRPRPRP$ going one way, or else $PLPLPLPL$ going the other.
III The Tonnetz as a Geometric Configuration
@@
—H. S. M. Coxeter (1950), Twelve Geometrical Essays
If we accept the widely-held view that Euler’s tonnetz is not a mere musico-mathematical curiosity but is fundamental to the structure of music (specifically, to that of the common practice period), then it makes sense to look at the mathematics of the tonnetz from every angle – group-theoretic, number-theoretic, set-theoretic, combinatoric and geometric – for insights that might be gained.
@@
Configurations of the form $\{n_{3}\}$ have attracted much attention and have been investigated at length. The inequality (1) shows that no $\{n_{3}\}$ configurations exist for $n\leq 6$. We draw on this fact a little later. It can be interesting to take stock of some of the configurations that exist for higher values of $n$.
@@
There is a unique configuration of type $\{7_{3}\}$, which can be modelled by the combinatoric geometry of the seven points and the seven lines of the projective plane based on the finite ‘binary’ field $\mathbb{Z}_{2}$. This is the so-called Fano plane. There is a unique combinatorial configuration of type $\{8_{3}\}$, which can be realized (i.e., modelled explicitly) by a system of eight points and eight lines in the complex projective plane $\mathbb{CP}^{2}$. No real coordinatization of $\{8_{3}\}$ on the Euclidean plane exists. The situation is a little more complicated for higher values of $n$. There are three distinct combinatorial configurations $\{9_{3}\}$ , all of which can be modelled as systems of points and lines in $\mathbb{R}^{2}$. One of these is the Pappus configuration mentioned earlier. All three of the configurations $\{9_{3}\}$ can be modelled over the rational numbers.
There are ten distinct combinatorial configurations $\{10_{3}\}$, nine of which can be modelled as geometric configurations in $\mathbb{R}^{2}$ and are rationally realizable. Among these is the Desargues configuration $\{10_{3}\}$, about which we have more to say later. But one of the combinatorial configurations $\{10_{3}\}$ does not admit a representation as a configuration of points and lines in $\mathbb{R}^{2}$. The realizability of a combinatorial configuration as a geometric configuration in $\mathbb{R}^{2}$ is hence a nontrivial property that may or may not hold.
The 31 distinct $\{11_{3}\}$ configurations were identified by Daublebsky von Sterneck (1894), all of which can be realized as geometric configurations in $\mathbb{R}^{2}$. The situation with $\{12_{3}\}$ is not completely straightforward. Daublebsky von Sterneck (1895) established the existence of 228 distinct configurations of this type. This number was accepted as definitive as recently as the work of Sturmfels & White (1990). But later Gropp (1993, 1997) showed that one case was missing from Daublebsky von Sterneck’s list: hence, in fact, the number of $\{12_{3}\}$ combinatorial configurations is 229.
Sturmfels & White (1990) were investigating the conjecture of Grünbaum that every $\{n_{3}\}$ that can be realized over the reals can be realized over the rationals. They established by use of diophantine computations that all the $\{11_{3}\}$ and $\{12_{3}\}$ cases considered by Daublebsky von Sterneck were realizable over $\mathbb{Q}$. Later it was recognized that the remaining configuration is also realizable over the rationals.
It follows by inspection of Figure 2 that this graph is the Levi graph of a combinatorial configuration of type $\{12_{3}\}$. In fact, matters can be taken further: the combinatorial configuration of the tonnetz can be realized as a geometric configuration of type $\{12_{3}\}$ in $\mathbb{R}^{2}$. This follows from a long line of work by geometers and combinatorialists (Daublebsky von Sterneck 1896, Sturmfels & White 1990, Gropp 1993,1997, Betten, Brinkmann & Pisanskic 2000, Alazemi & Betten 2014, and others) who have established that all 229 of the combinatorial $\{12_{3}\}$ configurations admit geometric realizations in $\mathbb{R}^{2}$ and indeed in the rational plane. That implies, as a consequence, that the tonnetz can be drawn as a configuration of points and lines in the Euclidian plane, the major chords being represented by points and the minor chords by lines, so that the points lie on lines in threes and the lines go through points in threes, in the relations dictated by the structure of the tonnetz.
@@
But which configuration? Once one gets this far, the problem that remains is to identify which one of 229 known configurations of type $\{12_{3}\}$ is the Eulerian tonnetz. We were able to identify it as the configuration D222. This remarkable configuration of points and lines, labelled with major and minor chords, which we call the Eulerian tonnetz configuration, is depicted in Figure 4. Thus, we have obtain a truly interesting and unexpected link between music and geometry! Summing up, we have the following:
@@
The Eulerian tonnetz is isomorphic to the self-dual configuration of type $\{12_{3}\}$ in $\mathbb{R}^{2}$ known as the D222 of Daublebsky von Sterneck. The twelve major chords are represented by points and the twelve minor chords are represented by lines. The thirty-six incidence relations between the twelve points and the twelve lines determine the edges of the associated Levi graph.
Figure 4: The Eulerian tonnetz admits a realization as a configuration on the Euclidean plane as a system of points and lines with a $\{12_{3}\}$ incidence scheme of Daublebsky von Sterneck type D222. Major chords are represented by points and minor chords by lines. Through each major vertex one finds the three minor lines associated with it; and on each minor line one finds the three associated major vertices. The $3p$-hexacycles such as $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$ form a closed sequence of four triangles, which can be easily identified, with the property that the vertices of each such triangle are inscribed in the lines of the next. The remaining twelve triangles of the configuration, which correspond to 2pp-hexacycles, can also be readily seen and counted.
@@ IV Cyclic Structure as a Basis for Musical Analysis
Clearly, any geometric configuration gives rise to a combinatorial configuration. It follows that any geometric configuration gives rise to a Levi graph, the white vertices corresponding to points and the black vertices to lines. Thus a triangle in $\mathbb{R}^{2}$ determines a hexagonal Levi graph. In particular, the hexacycles of the tonnetz are represented by triangles of the tonnetz configuration in $\mathbb{R}^{2}$. The 3pp-hexacycle containing the chord $CM$, for example, can be represented by a triangle in which the three major chords $CM$, EMEM, $A_{\flat}M$ correspond to points in $\mathbb{RP}^{2}$ and the three minor chords $Em$, $C_{\sharp}m$, $Cm$ correspond to the lines that join them. But the 2pp-hexacycles also correspond to triangles in the tonnetz configuration. For example, the 2pp-hexacycle $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$ is the triangle with vertices $CM,FM,A_{\flat}M$. It is then a simple visual exercise to count the total number of triangles in Figure 4, allowing us to determine altogether exactly sixteen different hexacycles in the tonnetz. The same exercise can be carried out in the Levi graph, since its structure is isomorphic to that of the configuration, but the counting is not so easy; and if one attempts this exercise in the setting of the traditional chicken-wire representation of the tonnetz, such as that shown in Figure 1, one can easily get lost, running about like a headless chicken. The point is that each of these representations of the tonnetz – tessellation, Levi graph, configuration – has merit in revealing aspects of the structure of the tonnetz.
Similarly, one sees that the octacycles of the tonnetz are represented by quadrilaterals of the tonnetz configuration. The 4pp-octacycle $\langle Cm$, $E_{\flat}M$, $E_{\flat}m$, $G_{\flat}M$, $F_{\sharp}m$ AMAM, AmAm, $CM\rangle$ arising here as a blade-like quadrilateral with vertices $E_{\flat}M,G_{\flat}M,AM,CM$ may look familiar since it is identical to the chord sequence of the ‘four-cornered hat’ starting at $Cm$ in Figure 2. But it is also identical to the chord sequence arising in Schubert’s Overture to Zauberharfe in the analysis of Cohn (2012) at pages 85-89. The symmetric beauty of Schubert’s composition seems to be associated with the fact that it is based on a 4pp-octacycle. In particular, the three 4pp-octacycles, which correspond to a system of three mutually exclusive blade-like ‘box-cutter’ quadrilaterals in the configuration, partition the tonnetz and Schubert’s trajectory in the Overture to Zauberharfe is confined to one of these sequences.
@@
It should be evident that any self-dual configuration is balanced. The triangle is thus a $\{3_{2}\}$ and for each $m\geq 3$ we can construct a self-dual configuration $\{m_{2}\}$ in the plane. On the other hand, a balanced configuration is not necessarily self-dual. This is not so obvious, since the combinatorial configurations associated with low order symmetric Levi graphs all turn out to be self-dual. In particular, among the balanced combinatorial configurations, the Fano $\{7_{3}\}$, the Möbius $\{8_{3}\}$, the Pappus $\{9_{3}\}$, the Desargues $\{10_{3}\}$, along with the other two $\{9_{3}\}$ configurations and the nine other $\{10_{3}\}$ configurations are all self-dual. But of the 31 configurations of type $\{11_{3}\}$, only 25 are self-dual, and of the 229 configurations of type $\{12_{3}\}$, only 95 are self-dual (Betten, Brinkmann & Pisanskic 2000). As we have remarked, all the $229$ combinatorial configurations of type $\{12_{3}\}$ can be represented in the real plane and indeed the rational plane. This applies in particular to the D222 (Alazemi & Betten 2014, Betten, Sturmfels & White 1997, Daublebsky von Sterneck 1895, 1903).
@@
That the twelve $2p$-hexacycles overlap in this way adds a depth of complexity to the scheme of triadic relations that may at first glance seem formidable but on reflection adds interest in a way that contrasts remarkably with the mutual disjointedness of the four $3p$-hexatonic cycles, hence allowing for architectural modalities. As an example, we mention the possibility of ‘surgical insertion’ of one cycle within another. Suppose we consider the 2pp-hexacycle (3) alongside the transposed cycle lying a fourth above,
@@
We observe also that beginning at the chord $CM$, the progression moving clockwise around the graph involves a sequence of ‘elevations’. More precisely, from the initial chord $CM^{(0)}=[C,E,G]$ we move to $Am^{(2)}=[C,E,A]$ by raising the $G$ to an $A$. Then we convert $Am^{(2)}$ to $Am^{(0)}$ by an action of the permutation group and perform another elevation, taking us from $[A,C,E]$ to $[A,C,F]$ by raising the note $E$ to an $F$. Then, converting the $FM^{(2)}$ to $FM^{(0)}$ by use of the permutation group and performing another elevation by raising the $C$ to a $D$, we get to $Dm^{(2)}$. In this way, we march from vertex to vertex through all twenty-four major and minor chords, giving us a Hamiltonian cycle (Harary 1969, pp. 65-70), i.e., a path that includes every vertex of the graph without repetitions. As Tymoczko (2012) remarks, a Hamiltonian path touches on all the vertices in a graph without passing through any of them twice, and is in that sense a generalization of the twelve-tone row.
@@
As usual with matters of artistic creativity, there are no fixed recipes for the modes in which the compositional elements should be employed; rather, merely suggestions of possibility. Nonetheless, it makes sense to let mathematics lead the way. On the matter of combining cycles together to produce new cycles, some clarity can be added if we recall the following well-known result from graph theory. Let $C_{1}$ and $C_{2}$ be cycles of a graph $G$ and suppose that the two cycles share an edge $e$ in common. Then there exists a cycle $C_{3}\subseteq\{C_{1}\cup C_{2}\}\backslash e$. That was indeed the setup, for example, with the hexacycles (3) and (6), which shared the edge $[FM,Fm]$, hence leading to a decacycle. In more detail, taking the union of the sets of the edges in the two hexacycles and removing the common edge, we were left with a set whose elements form a new cycle, a decacycle. Such operations can be pursued at greater length.  [^11] The idea of a cycle can be generalized by introducing the concept of a ‘cycle vector’ (Harary 1969, chapter 4). By a cycle vector in a graph $G$ we mean a collection of cycles of $G$ that are mutually edge-disjoint – that is to say, their edges do not overlap. A cycle vector is a formal sum of edge-disjoint cycles. The collection of all cycle vectors on a connected graph $G$ forms a vector space $\mathbb{V}$ over the binary field $\mathbb{F}(2)=\{0,1\}$ with the usual multiplication laws and the addition laws 0+0=0,  0+1=1,  1+1=00+0=0,\,\,0+1=1,\,\,1+1=0.
@@
The octatonic tonnetz cycles are of interest in their own right as objects of musical analysis on account of their relationship to so-called octatonic scales. An octatonic scale is defined to be a scale with eight notes in it, as opposed to the seven notes of the conventional major and minor scales, with the property that adjacent notes within the scale differ either by a semitone or a tone, alternating the gap at each step. One construction of such scales is as follows. There are three diminished seventh chords, each composed of four notes. If we form the union of two of these sets of four notes, this can be done in three ways, leading to three distinct sets of eight notes. Once these notes are ordered by pitch we obtain three distinct octatonic scales. The three diminished seventh chords,
$X=[C,E_{\flat},G_{\flat},A],\hskip 5.69046ptY=[D_{\flat},E,G,B_{\flat}],\hskip 5.69046ptZ=[D,F,A_{\flat},B]$ (8)
are obviously disjoint as sets. Taking the three distinct unions of pairs we obtain the three octatonic scales,
$\hskip 8.5359ptO_{12}=X\cup Y=[C,D_{\flat},E_{\flat},E,G_{\flat},G,A,B_{\flat}],$ (9)
$O_{23}=Y\cup Z=[D_{\flat},D,E,F,G,A_{\flat},B_{\flat},B],$ (10)
$O_{31}=Z\cup X=[D,E_{\flat},F,G_{\flat},A_{\flat},A,B,C].$ (11)
Any octatonic scale, starting at any chosen note, belongs to one of these three types, possibly after a cyclic permutation of its elements. Octatonic scales have been used to good effect by a number of well-known composers, among them Rimsky-Korsakov, Scriabin, Ravel, Stravinsky and Messiaen – and the use of these scales can be found in the works of earlier musicians as well (Messiaen 1956, Dallin 1964).

### v5

@@ II The Tonnetz as a Levi Graph
An $L$-transformation sends a major chord one notch counterclockwise on the Levi graph to the previous minor chord, and it sends a minor chord one step clockwise to the next major chord. An $R$-transformation sends a major chord notch clockwise on the Levi graph to the next minor chord, and it sends a minor chord one step counterclockwise to the previous major chord. A $P$-transformation sends any chord across the Levi graph to its mode reverse.
The general pattern of a $2p$-hexacycle is as follows. Beginning at the chord $CM$ the cycle moves two steps clockwise along the graph to $FM$; then it crosses the graph with a $P$-transformation to $Fm$; then it moves two steps counter-clockwise to $Cm$, and finally back to $CM$ with another $P$-transformation. From left to right the string of operations applied to the initial chord $CM$ in this example going clockwise thus takes the form $RLPRLP$, or else $PLRPLR$ going counterclockwise.
The other remaining eleven $2p$-hexacycles can be obtained by moving the initial point around in a circle of fourths, that is to say, from $CM$ to $FM$ to $B_{\flat}M$ and so on. Each tone is contained in three distinct $2p$-hexacycles. The $3p$-hexacycle (‘blunted triangle’) beginning at $CM$ is generated by the string $PLPLPL$ clockwise and the string $LPLPLP$ counterclockwise. This is the unique hexatonic cycle containing that chord. The $1p$-octacycle (‘octabeanie’) beginning at $CM$ is generated by the string $RLRLRLRP$. And the $4p$-octacycle (‘four-cornered hat’) beginning at $CM$ is generated by $RPRPRPRP$ going one way, or else $PRPRPRPR$ going the other.
III The Tonnetz as a Configuration
@@
—H. S. M. Coxeter, Twelve Geometrical Essays
If we accept the view held by many that Euler’s tonnetz is not a mere musico-mathematical curiosity but is fundamental to the structure of music (specifically, to that of the common practice period), then it makes sense to look at the mathematics of the tonnetz from every angle – group-theoretic, number-theoretic, set-theoretic, combinatoric and geometric – for insights that might be gained.
@@
Configurations of the form $\{n_{3}\}$ have attracted much attention and have been investigated at length. The inequality (1) shows that no $\{n_{3}\}$ configurations exist for $n\leq 6$. We draw on this fact a little later. It can be interesting to take stock of some of the configurations that exist for higher values of $n$. There is a unique configuration $\{7_{3}\}$, modelled by the seven points and seven lines of the finite projective plane based on the ‘binary’ field $\mathbb{Z}_{2}$. This is the Fano plane. There is a unique combinatorial configuration of type $\{8_{3}\}$, which can be realized explicitly by a system of eight points and eight complex projective lines in the complex projective plane $\mathbb{CP}^{2}$. No real coordinatization of $\{8_{3}\}$ on the Euclidean plane exists. There are three distinct combinatorial configurations $\{9_{3}\}$, all of which can be modelled as systems of points and lines in $\mathbb{R}^{2}$. One of these is the Pappus configuration mentioned earlier. All three of the configurations $\{9_{3}\}$ can be modelled over the rational numbers. There are ten distinct combinatorial configurations $\{10_{3}\}$, nine of which can be modelled as geometric configurations in $\mathbb{R}^{2}$ and are rationally realizable. Among these is the famous Desargues configuration $\{10_{3}\}$, about which we shall have more to say later. But one of the combinatorial configurations $\{10_{3}\}$ does not admit a representation as a geometric configuration over any field.
The 31 distinct configurations of type $\{11_{3}\}$ were identified by Daublebsky von Sterneck (1894), all of which can be realized in $\mathbb{R}^{2}$.  [^13] In his paper on the configurations $\{11_{3}\}$, Daublebsky von Sterneck (1894) says the following, which we can presume applies equally to his later work on configurations $\{12_{3}\}$: ‘Es handelt sich hier nur darum, alle möglichen Schemata, welche obige Eigenschaften besitzen, herzustellen und ist die Aufgabe insoferne ein Problem der Combinationslehre. Die im Folgenden verwendeten geometrischen Betrachtungen dienen bloss dazu, die combinatorischen Überlegungen zu unterstützen und zu erleichtern.’ In English: ‘Here the only goal is to create all possible systems which possess the above $\{11_{3}\}\,]$ properties and in this respect the task is a problem of combinatorics. The geometric considerations used in the following serve merely to support and clarify the combinatorial considerations.’ The remark is interesting in the present context, since one can take the view that the tonnetz is primarily a combinatoric object, well-represented as such by its Levi graph, and that its geometric representation as a configuration in $\mathbb{R}^{2}$ mainly serves to support and clarify the combinatorial considerations. This is the minimal position that one can take with reference to the role of combinatorial configurations in music. Nonetheless, combinatorics and geometry have come a long way since the days of Daublebsky von Sterneck and from a modern perspective we take a more unified approach to the relations between these disciplines – and hence the view that music is somehow deeply entangled with this unity. The test as to whether this is really the case is to see whether entertaining such a view is fruitful in the new ideas that it generates in the context of mathematical music theory. The situation with $\{12_{3}\}$ is not completely straightforward. Daublebsky von Sterneck (1895, 1903) established the existence of 228 distinct configurations of this type. This number was accepted as definitive as recently as the work of Sturmfels & White (1990). Then Gropp (1991, 1993, 1997) showed that one case was missing from Daublebsky von Sterneck’s list: hence, in fact, the number of $\{12_{3}\}$ combinatorial configurations is 229. Sturmfels & White (1990) were investigating the conjecture of Grünbaum that every $\{n_{3}\}$ that can be realized over the reals can be realized over the rationals. They established by use of diophantine computations that all the $\{11_{3}\}$ and $\{12_{3}\}$ cases considered by Daublebsky von Sterneck were realizable over $\mathbb{Q}$. Thus it follows from a long line of work by geometers and combinatorialists spanning more than a century that all 229 of the combinatorial $\{12_{3}\}$ configurations admit geometric realizations in $\mathbb{R}^{2}$ and indeed in the rational plane.  [^14] See Betten, Brinkmann & Pisanskic (2000), Gropp (2004), Grünbaum (2009) and Alazemi & Betten (2014) for details of the history of the subject and for tables and constructions for $\{12_{3}\}$ and $\{n_{3}\}$ for higher values of $n$.
Now we are in a position to say how the tonnetz fits into this picture. We know by inspection of Figure 2 that the tonnetz can be represented as the Levi graph of a combinatorial configuration of type $\{12_{3}\}$. But we also know that each such combinatorial configuration $\{12_{3}\}$ can be realized as a geometric configuration $\{12_{3}\}$ in $\mathbb{R}^{2}$. That implies, as a consequence, that the tonnetz is such a configuration – that it can be drawn as a geometric configuration of points and lines in the Euclidean plane, the major chords being represented by points and the minor chords by lines, so that the points lie on lines in threes and the lines go through points in threes.
@@
But which configuration? Once one gets this far, the problem that remains is to determine which one of 229 known configurations of type $\{12_{3}\}$ is the Eulerian tonnetz. We have been able to identify it as the configuration D222. This remarkable configuration of points and lines, labelled with major and minor chords, which we call the tonnetz configuration, is depicted in Figure 4, in line with the original drawing of Daublebsky von Sterneck (1895). Thus, we have obtained a truly interesting and unexpected link between music and geometry! Summing up, we have the following:
@@
The Eulerian tonnetz can be represented in $\mathbb{R}^{2}$ as a self-dual configuration $\{12_{3}\}$ of Daublebsky von Sterneck type D222. The twelve major chords are represented by points and the twelve minor chords are represented by lines. Three of the points lie on each line and three of the lines pass through each point. The thirty-six incidence relations between the twelve points and the twelve lines determine the edges of the associated Levi graph.
Figure 4: The Eulerian tonnetz admits a realization as a geometric configuration on the Euclidean plane with a $\{12_{3}\}$ incidence scheme of Daublebsky von Sterneck type D222. Major chords are represented by points and minor chords by lines. Through each point representing a major chord one finds three lines representing the minor chords associated with it; and on each such line representing a minor chord one finds three points representing the three associated major chords. The four $3p$-hexacycles such as $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$ form a closed sequence of four triangles, which can be easily identified, with the property that the vertices of each such triangle are inscribed in the lines of the next. The remaining twelve triangles of the configuration, which correspond to 2pp-hexacycles, can also be readily seen and counted.
@@ IV Cyclic Structure as a Basis for Musical Analysis
Clearly, any geometric configuration gives rise to a combinatorial configuration. It follows that any geometric configuration gives rise to a Levi graph, the white vertices corresponding to points and the black vertices to lines. Thus a triangle in $\mathbb{R}^{2}$ determines a hexagonal Levi graph. In particular, the hexacycles of the tonnetz are represented by triangles of the tonnetz configuration in $\mathbb{R}^{2}$. The 3pp-hexacycle containing the chord $CM$, for example, can be represented by a triangle in which the three major chords $CM$, EMEM, $A_{\flat}M$ correspond to points in $\mathbb{RP}^{2}$ and the three minor chords $Em$, $C_{\sharp}m$, $Cm$ correspond to the lines that join them. But the 2pp-hexacycles also correspond to triangles in the tonnetz configuration. For example, the 2pp-hexacycle $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$ is the triangle with vertices $CM,FM,A_{\flat}M$. It is then a simple visual exercise to count the total number of triangles in Figure 4, allowing us to determine altogether exactly sixteen different hexacycles in the tonnetz. The same exercise can be carried out in the Levi graph, since its structure is isomorphic to that of the configuration, but the counting is not so easy; and if one attempts this exercise in the setting of the traditional chicken-wire representation of the tonnetz, such as that shown in Figure 1, one can easily get lost, running about like a headless chicken. The point is that each of these representations of the tonnetz – tessellation, Levi graph, configuration – has merit in revealing aspects of the structure of the tonnetz. Here we have a good example of how the geometric structure of the tonnetz as a configuration in $\mathbb{R}^{2}$ makes is easy to identify certain key elements of musical structure – namely, the hexacycles – which are but the humble triangles of the configuration.
Similarly, one sees that the octacycles of the tonnetz are represented by quadrilaterals of the tonnetz configuration. The 4pp-octacycle $\langle Cm$, $E_{\flat}M$, $E_{\flat}m$, $G_{\flat}M$, $F_{\sharp}m$ AMAM, AmAm, $CM\rangle$ arising here as a blade-like quadrilateral with vertices $E_{\flat}M,G_{\flat}M,AM,CM$ may look familiar. It is identical to the chord sequence of the ‘four-cornered hat’ starting at $Cm$ in Figure 2. But it is also identical to the chord sequence arising in Schubert’s Overture to Zauberharfe (see Cohn 2012, pp. 85-89). The symmetric beauty of Schubert’s composition seems to be associated with the fact that it is based on a 4pp-octacycle. In particular, the three 4pp-octacycles, which correspond to a system of three mutually exclusive blade-like ‘box-cutter’ quadrilaterals in the configuration, partition the tonnetz and Schubert’s trajectory in the Overture to Zauberharfe is confined to one of these sequences.  [^15] There is no ‘musical information’ in a combinatorial configuration that is not already implicit in the associated Levi graph. This is because they are isomorphic. Nonetheless, the visualization of geometric configurations makes them ideal for the representation of certain musical ideas. A good example of this is the surprising visualization of the octacyclic progression of dominant sevenths and half-diminished sevenths in the Tristan tonnetz presented in Section X in our discussion of the final scene of Götterdämmerung.
@@
It should be evident that any self-dual configuration is balanced. The triangle is thus a $\{3_{2}\}$ and for each $m\geq 3$ we can construct a self-dual configuration $\{m_{2}\}$ in the plane. On the other hand, a balanced configuration is not necessarily self-dual. This is not so obvious, since the combinatorial configurations associated with low order symmetric Levi graphs all turn out to be self-dual. In particular, among the balanced combinatorial configurations, the Fano $\{7_{3}\}$, the Möbius $\{8_{3}\}$, the Pappus $\{9_{3}\}$, the Desargues $\{10_{3}\}$, along with the other two $\{9_{3}\}$ configurations and the nine other $\{10_{3}\}$ configurations are all self-dual. But of the 31 configurations of type $\{11_{3}\}$, only 25 are self-dual, and of the 229 configurations of type $\{12_{3}\}$, only 95 are self-dual (Betten, Brinkmann & Pisanskic 2000).
As we have remarked, all the $229$ combinatorial configurations of type $\{12_{3}\}$ can be represented in the real plane and indeed the rational plane. This applies in particular to the D222 (Daublebsky von Sterneck 1895, 1903, Sturmfels & White 1990, Betten, Brinkmann & Pisanskic 2000, Alazemi & Betten 2014).
@@
That the twelve $2p$-hexacycles overlap in this way adds a depth of complexity to the scheme of triadic relations that may at first glance seem formidable but on reflection adds interest in a way that contrasts remarkably with the mutual disjointedness of the four $3p$-hexacycles, hence allowing for architectural modalities. As an example, we mention the possibility of ‘surgical insertion’ of one cycle within another. Suppose we consider the 2pp-hexacycle (3) alongside the transposed cycle lying a fourth above,
@@
We observe also that beginning at the chord $CM$ in Figure 2, the progression moving clockwise around the perimeter of the Levi graph involves a sequence of transformations that we call ‘elevations’. More precisely, from the initial chord $CM^{(0)}=[C,E,G]$ we move to $Am^{(2)}=[C,E,A]$ by raising the $G$ to an $A$. Then we convert $Am^{(2)}$ to $Am^{(0)}$ by an action of the permutation group and perform another elevation, taking us from $[A,C,E]$ to $[A,C,F]$ by raising the note $E$ to an $F$. Then, converting the $FM^{(2)}$ to $FM^{(0)}$ by use of the permutation group and performing another elevation by raising the $C$ to a $D$, we get to $Dm^{(2)}$. In this way, we march from vertex to vertex through all twenty-four major and minor chords, giving us a Hamiltonian cycle (Harary 1969, pp. 65-70), i.e., a path that includes every vertex of the graph without repetitions. As Tymoczko (2012) remarks, a Hamiltonian path touches on all the vertices in a graph without passing through any of them twice, and is in that sense a generalization of the idea of a twelve-tone row.  [^16] See Tymoczko (2011b, 2020) on geometrical models of voice leading.
@@
As usual with matters of artistic creativity, there are no fixed recipes for the modes in which the compositional elements should be employed; rather, merely suggestions of possibility. Nonetheless, it makes sense to let mathematics lead the way. On the matter of combining cycles together to produce new cycles, some clarity can be added if we recall the following well-known result from graph theory. Let $C_{1}$ and $C_{2}$ be cycles of a graph $G$ and suppose that the two cycles share an edge $e$ in common. Then there exists a cycle $C_{3}\subseteq\{C_{1}\cup C_{2}\}\backslash e$. That was indeed the setup, for example, with the hexacycles (3) and (6), which shared the edge $[FM,Fm]$, hence leading to a decacycle. In more detail, taking the union of the sets of the edges in the two hexacycles and removing the common edge, we were left with a set whose elements form a new cycle, a decacycle. Such operations can be pursued at greater length.  [^17] See Cohn (2012) for discussions of the cycles of the tonnetz. The idea of a cycle can be generalized by introducing the concept of a ‘cycle vector’ (Harary 1969, chapter 4). By a cycle vector in a graph $G$ we mean a collection of cycles of $G$ that are mutually edge-disjoint – that is to say, their edges do not overlap. A cycle vector is a formal sum of edge-disjoint cycles. The collection of all cycle vectors on a connected graph $G$ forms a vector space $\mathbb{V}$ over the binary field $\mathbb{F}(2)=\{0,1\}$ with the usual multiplication laws and the addition laws 0+0=0,  0+1=1,  1+1=00+0=0,\,\,0+1=1,\,\,1+1=0.
@@
The $4p$-octacycles are of interest in their own right as objects of musical analysis on account of their relationship to octatonic scales. An octatonic scale is defined to be a scale with eight tones in it, as opposed to the seven tones of the conventional major and minor scales, with the property that adjacent notes within the scale differ either by a semitone or a tone, alternating the gap at each step. One construction of such scales is as follows. There are three diminished seventh chords, each composed of four notes. If we form the union of two of these sets of four notes, this can be done in three ways, leading to three distinct sets of eight notes. Once these notes are ordered by pitch we obtain three distinct octatonic scales. The three diminished seventh chords, $X=[C,E_{\flat},G_{\flat},A]$, $Y=[D_{\flat},E,G,B_{\flat}]$, and $Z=[D,F,A_{\flat},B]$ are disjoint as sets. Taking three distinct unions of pairs we get the three octatonic scales, $O_{12}=X\cup Y=[C,D_{\flat},E_{\flat},E,G_{\flat},G,A,B_{\flat},C]$, $O_{23}=Y\cup Z=[D_{\flat},D,E,F,G,A_{\flat},B_{\flat},B,D_{\flat}]$ and $O_{31}=Z\cup X=[D,E_{\flat},F,G_{\flat},A_{\flat},A,B,C,D]$. Any octatonic scale, starting at any chosen note, belongs to one of these three types, possibly after a cyclic permutation of its elements. Octatonic scales have been used to good effect by a number of well-known composers, among them Rimsky-Korsakov, Scriabin, Ravel, Stravinsky and Messiaen – and the use of these scales can be found in the works of earlier musicians as well (Messiaen 1956, Dallin 1964).